Guy Wire. How long is a guy wire reaching from the top of a 9 ft pole to a point on the ground 6 feet from the pole. What is the exact length in feet?
Answer by rfer(16322)

a^2+b^2=c^2
81+36=117
sqrt 117=10.81665383 ft